OpenStudy (anonymous):

A person X is running around a circular track completing one round every 40 seconds. Another person Y running in the opposite direction meets X every 15 second. The time, expressed in seconds, taken by Y to complete one round is?

OpenStudy (anonymous):

@hartnn

OpenStudy (tkhunny):

If person X can complete the track in 40 seconds, and If person X and Person Y, working together, can complete the track in 15 seconds, How long does it take person Y to complete the track alone?

OpenStudy (anonymous):

x=40 x+y=15 y=15-x y=15-40 -25 seconds :)

OpenStudy (tkhunny):

That's no good. Perhaps you have seen this problem: If Billy can paint a doghouse in 40 minutes, and If Billy and Wanda, working together, can paint the doghouse in 15 minutes, How long does it take Wanda to paint the doghouse alone?

OpenStudy (anonymous):

don't know..what is wrong in the method

OpenStudy (tkhunny):

It's silly. How can it be -25? No one runs a lap in -25 seconds! That makes absolutely no sense. This is why you should do a little reasonableness before you start. Okay, it has to be positive. If I get anything negative, something went horribly wrong. Okay, if they were going the same speed, they would meet half way around every time, or 20 seconds. So, Person Y must be FASTER - completing a circuit in less than 40 seconds. There, some reasonableness. The answer MUST be between 0 and 40. Anything less than zero or greater than 40 is no good and we'll have to rethink. How about that paint problem. SURELY you have worked one of those. No?

OpenStudy (anonymous):

nope i dont remember..

OpenStudy (tkhunny):

Well, then you cannot solve this problem. What's your next plan?

OpenStudy (tkhunny):

X: 40 seconds per track X: 1/40 th of the track in 1 second Y: M seconds per track Y: 1/M th of the track in 1 second Together: 15 seconds per track Together: 1/15 th of the track in 1 second 1/40 + 1/M = 1/15 Try that.